# Extracting memorized pieces of (copyrighted) books from open-weight language models

### v3

@@ 8.2 It may be practical to generate infringing copies in some cases
Traditional extraction measurements vs. reconstruction of (whole) books.  Reliable extraction of large segments of text would require other approaches, like the reconstruction procedure that we apply to Harry Potter and the Sorcerer’s Stone with Llama 3.1 70B (Section 7). This result is an extreme case and is not representative of most of the other books we have tested (Appendix F.2). Nevertheless, it is significant: in contrast to claims by some defendants [27], it illustrates that, for LLMs that exhibit high degrees of memorization of specific works, it may be possible to generate copies or derivative works of those works. General assessments of the extent of this type of behavior, and how easy it is to elicit it, require more detailed exploration.
Practicality of extraction and responsibility for infringement.  Even though our results suggest it may be possible to reconstruct large segments of copyrighted text from an LLM, they do not imply that it will always be easy to do so. This may affect who courts view as directly liable for any output infringement [76]. As a policy matter, the case for liability against the model developer is stronger if simple prompts generate infringing outputs. The case is weaker if a user (e.g., a plaintiffs’ lawyer) employs more intensive strategies to try to generate infringing material. In that case, the adversarial user—who perhaps runs hundreds or thousands of prompts to try to get one output that is infringing—seems more like a direct volitional actor misusing the model [32]; they should likely bear some responsibility for the infringement. [^17] Models are used at scale—in some cases, by millions of users. Hypothetically, a million users interacting with a particular model might independently prompt the model with the prefix of a memorized sequence. Even for a verbatim suffix that (in expectation) appears only once for every 1,0001{,}000 prompts (i.e., $p_{\bm{z}}{=}\tau_{\text{min}}{=}0.1\%$), in expectation 1,0001{,}000 of these users could expect to see the verbatim memorized suffix in the output. By contrast, in several cases (e.g., the first Harry Potter book, 1984), it does not take a thousand or even hundreds of generations to get large amounts of text; it takes only a handful. And in one case that we have identified so far (Harry Potter and the Sorcerer’s Stone), it is possible to generate the book directly from a short starting prompt and boilerplate generation code (Section 7). We will address this further in future work.
@@
Nonetheless, the fact that memorization occurs at all creates a point of distinction from Authors Guild v. Google [6]. In this case, the court held that Google’s scanning of books to create a searchable database, and Google’s public display of limited text snippets on Google Books, was fair use. The court also considered and rejected an argument that hackers could gain unauthorized access to Google’s internal books database that supported the service, and thereby obtain plaintiffs’ entire works [1, at 227-28]. In contrast, our results show that it is possible to recreate some of the content of some books by “hacking” the model itself. We think it unlikely as a practical matter that people will use the model in this way; there are easier and more effective ways to pirate a book. But at a minimum, it complicates the copyright fair use analysis.
@@ 8.3 Class actions are ill-suited for copyright claims about memorization
To be certified as a class action, plaintiffs generally have to show that all putative class members share common legal and factual issues, as well as suffer common injuries. Because many of the pending cases are proceeding as class actions, plaintiffs will therefore need to demonstrate that all book owners meet this standard so that a court can certify the class and treat their claims together [44].
As is clear from our results, this may be impossible. Basic questions about whether a particular model actually incorporated any significant expression from any individual plaintiff’s book cannot be generalized. Rather, it varies case by case: the extent of memorization differs widely from model to model, and even from work to work within a model (Section 6). This makes it difficult to assess, on a class-wide basis, whether a particular model copied a particular work and whether, for that model, infringing outputs based on memorization are even possible. Indeed, we show that memorization of works varies for many of the actual named authors in lawsuits, and even across books for the same plaintiffs (e.g., Ta-Nehisi Coates, see Figure 6; Appendices D.3.6D.3.8, & E.3). Answering questions about the degree to which a specific LLM has memorized a specific work requires experimental evidence—for example, running every book through tests similar to the ones we performed here. Some plaintiffs may be able to show copying, but others will not. Courts generally deny class certification in such circumstances [e.g., 131].
@@ 9 Conclusion
Building on recent work on probabilistic discoverable extraction [60] (Section 2), we develop (Section 4) and validate (Section 5) a sliding-window procedure to estimate the extent to which open-weight models memorize different books contained in the Books3 training dataset. Our results show that the extent of verbatim memorization of books in open-weight LLMs is more significant than previously described. They also show that memorization varies widely from model to model and from book to book within each model, as well as varying in different parts of individual books (Section 6). We find that, with respect to our sliding-window probabilistic extraction procedure, most models do not memorize most books. But this is not always the case. In extreme instances, a model may memorize a book so significantly that it is possible to generate the entire book near-verbatim, as we have done with Llama 3.1 70B for Harry Potter and the Sorcerer’s Stone (Section 7).
Altogether, our results complicate current disputes over copyright infringement, both by rejecting easy claims made by both sides about how models work and by demonstrating that there is no single answer to the question of how much a model memorizes (Sections 38). There is much more work to do. We only ran book-specific, verbatim-extraction experiments for $50$ books in Books3. In future work, we intend to test significantly more books and models, as well as possible explanations for memorization variability and for the high level of memorization of some specific works by particular models.
@@
The computing for this project was performed on the Sherlock cluster. We would like to thank Stanford University and the Stanford Research Computing Center for providing computational resources and support that contributed to these research results. We thank Jamie Hayes for chatting about additional interesting aspects concerning the $(n,p)$-discoverable extraction method [60] that we did not cover in that original paper. We thank Timothy B. Lee for his extremely clear explanation of our work [78], which helped us simplify the presentation of the math in this paper. We would additionally like to thank Alexandra Chouldechova, danah boyd, Danny Wilf-Townsend, Florian Tramèr, Franziska Boenisch, James Grimmelmann, Katherine Lee, Kevin Klyman, Luis Villa, Matthew Jagielski, Matthew Sag, Milad Nasr, Nicholas Carlini, Pamela Samuelson, Paul Cantor, Stella Biderman, and others for helpful questions, comments, and feedback about earlier versions of this project.
@@ A.1 Additional notes on the “careless people” example
The headline figure we include in Figure 1 (Section 2) is meant to be illustrative—an instance of, not a general statement about, memorization and extraction. This example also plays a role in some of our validity experiments (Section 5 and Appendix E). The text is drawn from The Great Gatsby [46]; it is one of the most famous quotes from the book. It is also a short, $57$-token sequence: the prefix is $25$ tokens and the suffix is $32$ (with the Llama 1 tokenizer). We do not focus on such short suffixes in our analysis, for reasons that are discussed both in the main text and in our experimental details below. While extraction of verbatim $32$-token suffixes is often accepted as valid evidence for memorization [9], it is now more common practice to use $50$-token suffixes. (Regarding validity, also refer to Section 5 and Appendix E.) We compute the extraction probability with respect to top-$k$ sampling ($T=1$, $k=40$). We do this by transforming the model logits according to the sampling parameters of this scheme. (See more about this below in Appendix A.3.) It is also the case that very famous quotes like this are duplicated in many locations in training-data sources (e.g., blogs, essays, personal web pages, etc.) We are not claiming (here, or in general) that this sequence can be extracted because a given book (here, The Great Gatsby) was included in Books3; it could be extracted (or more extractable) because it is duplicated in many places in the training data. Of course, one of those copies is in Books3, if the entirety of Books3 is included in the model’s training data. We discuss this further in Appendix E.
@@ A.2 Metrics
The metric that we use for memorization comes from Hayes et al. [60]. The authors provide a different framing for the metric. Instead of directly comparing $p_{\bm{z}}$ values—for the same sequence across LLMs, across multiple sequences for the same LLM, etc.—Hayes et al. [60] specify a threat model with an adversary that can query the model an arbitrary number of times with the same prompt—an adversary that has $n$ independent attempts to extract a sequence $\bm{z}$} with at least probability $p$. This is where their metric derives its name, $\bm{(n,p)}$-discoverable extraction:
Definition 1 ($(n,p)$-discoverable extraction, from Hayes et al. [60]).
@@
On its own, this type of method is well-suited to completion models that continue the prompt text, rather than instruction-tuned chatbots that reply in a conversational style. We therefore only study completion models and, for brevity, refer to these models by name, version, and size. Also note that this definition can be (at least in principle) easily adapted to non-verbatim extraction by testing if a generation is within $\epsilon$ distance from the target suffix for a given distance metric. Despite these differences in framing, in practice for verbatim extraction, computing this metric just involves calculating sequence probabilities $p_{{\bm{z}},\theta,\phi}$. For sequence $\bm{z}$}, this is the suffix probability with respect to an LLM’s base distribution $\theta$ and decoding scheme $\phi$ (e.g., greedy, top-$k$) that determines which tokens are considered at each step and how their logits are renormalized into conditional probabilities (Appendix A.3). That is,
@@
We will always note the particular LLM $\theta$ and decoding scheme $\phi$ being evaluated, and so, for brevity, will refer to this quantity as simply $p_{\bm{z}}$. In practice, we compute $p_{\bm{z}}$ as the $\exp$ of the sum of conditional $\log$ probabilities of the tokens in the sequence, as this is more numerically stable than multiplying together the conditional probabilities of the tokens in the sequence. (See Section 2.) In Figure 10 (left), we show examples of $p_{\bm{z}}$ for the “careless people” quote (split into prefix and suffix as in Figure 1) for different models. (This is Figure 2 in Section 2, with Pythia 12B added.) Once one has computed $p_{\bm{z}}$, they can use $p_{\bm{z}}$ to compute $n$ for a given $p$ (for $p_{\bm{z}}\in(0,1)$) with
@@
Since $p_{\bm{z}}$ is the probability of generating the suffix with $1$ prompt to the model, the probability of not generating it in $1$ prompt is $(1-p_{\bm{z}})$. The probability of not generating it in $n$ independent prompts to the model (with the same prompt) is $(1-p_{\bm{z}})^{n}$, and so the probability of generating it in $n$ prompts is $1$ minus this probability, i.e., $1-(1-p_{\bm{z}})^{n}$. As an intuition for this presentation of the metric, think about a fair coin flip coming up heads. Here, $p_{\bm{z}}=0.5$ (the probability of heads in one flip). With two actual flips ($n=2$), the probability of flipping heads at least once is $1-(1-0.5)^{2}=0.75=p$; when $n=10$, $p=0.9990$. We visualize this intuition in Figure 11 (left) with coin flips, and show how it translates to suffix generation probabilities for Llama 1 30B and the “careless people” quote, for which $p_{\bm{z}}=35.2\%$ (right). As is clear Equation (6), for a given $p_{\bm{z}}$, one can pick a probability threshold $p$ and get the corresponding number of prompts $n$ (or vice versa). Figure 10 (right) visualizes how $n$ changes for different settings of $p$ for the “careless people” quote. In this paper, we re-frame $(n,p)$-discoverable extraction to emphasize the quantity $p_{\bm{z}}$ (A.2) that we actually compute: the probability of extracting a sequence $\bm{z}$} (for the given LLM, hyperparameter-configured decoding scheme, and suffix start index location within sequence $\bm{z}$}). This is also useful for plotting distributions over $p_{\bm{z}}$ for a given book—to see how probabilities vary for sequences across a book, or for a given sequence across different models. We provide such comparisons in Section 6.In practice, for LLMs, we can compute $p_{\bm{z}}$ for verbatim extraction without prompting $n$ times. We compute sequence probabilities directly from the logits we obtain from running sequences through the LLM, rather than generating multiple suffixes for each prefix. (See Appendix A.3.) These two approaches are functionally equivalent for verbatim extraction: the computed sequence probabilities correspond to the statistically expected frequencies of verbatim extracted outputs we would observe if we were to generate a large number of sequences for the same prompt. Hayes et al. [60], the authors of this extraction approach, confirm this with ample experimental evidence. We do, as well, though omit these results for brevity because they follow directly from what $p_{\bm{z}}$ is measuring.We limit ourselves to verbatim extraction in this paper; it is significantly computationally cheaper to do so because we can operate directly on model logits with one forward pass. (See Section 2 and Appendix A.3 for more details.) As a result, our work here (in several respects) only scratches the surface: there are various different ways to instantiate this extraction methodology in practice, we only explore a limited set of open-weight models, and we only test $50$ of the nearly 200,000200{,}000 books in Books3.We further discuss how we determine extraction success for probabilistic extraction in Appendix C, where we describe how extraction success for different sequences can be aggregated into an overall extraction rate. In the main text, we discuss determining extraction success in Section 2.1.

### v4

@@ 8.2 It may be practical to generate infringing copies in some cases
Traditional extraction measurements vs. reconstruction of (whole) books.  Reliable extraction of large segments of text would require other approaches, like the reconstruction procedure that we apply to Harry Potter and the Sorcerer’s Stone with Llama 3.1 70B (Section 7). This result is an extreme case and is not representative of most of the other books we have tested (Appendix F.3). Nevertheless, it is significant: in contrast to claims by some defendants [concordanthropic], it illustrates that, for LLMs that exhibit high degrees of memorization of specific works, it may be possible to generate copies or derivative works of those works. General assessments of the extent of this type of behavior, and how easy it is to elicit it, require more detailed exploration.
Practicality of extraction and responsibility for infringement.  Even though our results suggest it may be possible to reconstruct large segments of copyrighted text from an LLM, they do not imply that it will always be easy to do so. This may affect who courts view as directly liable for any output infringement [lee2023talkin]. As a policy matter, the case for liability against the model developer is stronger if simple prompts generate infringing outputs. The case is weaker if a user (e.g., a plaintiffs’ lawyer) employs more intensive strategies to try to generate infringing material. In that case, the adversarial user—who perhaps runs hundreds or thousands of prompts to try to get one output that is infringing—seems more like a direct volitional actor misusing the model [costar]; they should likely bear some responsibility for the infringement. [^17] Models are used at scale—in some cases, by millions of users. Hypothetically, a million users interacting with a particular model might independently prompt the model with the prefix of a memorized sequence. Even for a verbatim suffix that (in expectation) appears only once for every 1,0001{,}000 prompts (i.e., $p_{\bm{z}}{=}\tau_{\text{min}}{=}0.1\%$), in expectation 1,0001{,}000 of these users could expect to see the verbatim memorized suffix in the output. By contrast, in several cases (e.g., the first Harry Potter book, 1984), it does not take a thousand or even hundreds of generations to get large amounts of text; it takes only a handful. And in one case that we have identified so far (Harry Potter and the Sorcerer’s Stone), it is possible to generate the book directly from a short starting prompt and boilerplate generation code (Section 7). We will address this further in future work.
@@
Nonetheless, the fact that memorization occurs at all creates a point of distinction from Authors Guild v. Google [authorsguild]. In this case, the court held that Google’s scanning of books to create a searchable database, and Google’s public display of limited text snippets on Google Books, was fair use. The court also considered and rejected an argument that hackers could gain unauthorized access to Google’s internal books database that supported the service, and thereby obtain plaintiffs’ entire works [804fd, at 227-28]. In contrast, our results show that it is possible to recreate some of the content of some books by “hacking” the model itself. We think it unlikely as a practical matter that people will use the model in this way; there are easier and more effective ways to pirate a book. But at a minimum, it complicates the copyright fair use analysis.
@@ 8.3 Class actions are ill-suited for copyright claims about memorization
To be certified as a class action, plaintiffs generally have to show that all putative class members share common legal and factual issues, as well as suffer common injuries. Because many of the pending cases are proceeding as class actions, plaintiffs will therefore need to demonstrate that all book owners meet this standard so that a court can certify the class and treat their claims together [class].
As is clear from our results, this may be impossible. Basic questions about whether a particular model actually incorporated any significant expression from any individual plaintiff’s book cannot be generalized. Rather, it varies case by case: the extent of memorization differs widely from model to model, and even from work to work within a model (Section 6). This makes it difficult to assess, on a class-wide basis, whether a particular model copied a particular work and whether, for that model, infringing outputs based on memorization are even possible. Indeed, we show that memorization of works varies for many of the actual named authors in lawsuits, and even across books for the same plaintiffs (e.g., Ta-Nehisi Coates, see Figure 6; Appendices D.3.6D.3.8, & E.3). Answering questions about the degree to which a specific LLM has memorized a specific work requires experimental evidence—for example, running every book through tests similar to the ones we performed here. Some plaintiffs may be able to show copying, but others will not. Courts generally deny class certification in such circumstances [e.g., walmart].
@@ 9 Conclusion
Building on recent work on probabilistic discoverable extraction [hayes2025measuringmemorizationlanguagemodels] (Section 2), we develop (Section 4) and validate (Section 5) a sliding-window procedure to estimate the extent to which open-weight models memorize different books contained in the Books3 training dataset. Our results show that the extent of verbatim memorization of books in open-weight LLMs is more significant than previously described. They also show that memorization varies widely from model to model and from book to book within each model, as well as varying in different parts of individual books (Section 6). We find that, with respect to our sliding-window probabilistic extraction procedure, most models do not memorize most books. But this is not always the case. In extreme instances, a model may memorize a book so significantly that it is possible to generate the entire book near-verbatim, as we have done with Llama 3.1 70B for Harry Potter and the Sorcerer’s Stone (Section 7).
Altogether, our results complicate current disputes over copyright infringement, both by rejecting easy claims made by both sides about how models work and by demonstrating that there is no single answer to the question of how much a model memorizes (Sections 3 & 8). There is much more work to do. We only ran book-specific, verbatim-extraction experiments for $50$ books in Books3. In future work, we intend to test significantly more books and models, as well as possible explanations for memorization variability and for the high level of memorization of some specific works by particular models.
@@
The computing for this project was performed on the Sherlock cluster. We would like to thank Stanford University and the Stanford Research Computing Center for providing computational resources and support that contributed to these research results. We thank Jamie Hayes for chatting about additional interesting aspects concerning the $(n,p)$-discoverable extraction method [hayes2025measuringmemorizationlanguagemodels] that we did not cover in that original paper. We thank Timothy B. Lee for his extremely clear explanation of our work [lee2025arstechnica], which helped us simplify the presentation of the math in this paper. We would additionally like to thank Alexandra Chouldechova, danah boyd, Danny Wilf-Townsend, Florian Tramèr, Franziska Boenisch, James Grimmelmann, Katherine Lee, Kevin Klyman, Luis Villa, Matthew Jagielski, Matthew Sag, Milad Nasr, Nicholas Carlini, Pamela Samuelson, Paul Cantor, Stella Biderman, and others for helpful questions, comments, and feedback about earlier versions of this project.
@@
  • A
    Additional notes on memorization and probabilistic extraction ........................................................................................................................................................................A
    • A.1
      Additional notes on the “careless people” example ........................................................................................................................................................................A.1
    • A.2
      Metrics ........................................................................................................................................................................A.2
    • A.3
      Computing sequence probabilities in one forward pass ........................................................................................................................................................................A.3
  • B
    Testing our measurement pipeline ........................................................................................................................................................................B
  • C
    Experiments on extraction rates ........................................................................................................................................................................C
    • C.1
      Setup ........................................................................................................................................................................C.1
    • C.2
      Experimental procedure ........................................................................................................................................................................C.2
    • C.3
      High-level takeaways ........................................................................................................................................................................C.3
  • D
    Sliding-window experiments with our probabilistic extraction procedure ........................................................................................................................................................................D
    • D.1
      Setup ........................................................................................................................................................................D.1
    • D.2
      Sliding-window probabilistic extraction procedure ........................................................................................................................................................................D.2
    • D.3
      Book-specific results ........................................................................................................................................................................D.3
    • D.4
      Estimating how much of a book is memorized by a model ........................................................................................................................................................................D.4
      • D.4.1
        Quantifying extraction coverage ........................................................................................................................................................................D.4.1
      • D.4.1
        Extended extraction coverage results........................................................................................................................................................................D.4.2
    • D.5
      Selection of extracted sequences ........................................................................................................................................................................D.5
  • E
    Measurement validity ........................................................................................................................................................................E
    • E.1
      Motivating our validity experiments ........................................................................................................................................................................E.1
      • E.1.1
        Setting $\tau_{\text{min}}$ and extending to models with unknown training data ........................................................................................................................................................................E.1.1
      • E.1.2
        Relating greedy-decoded and probabilistic discoverable extraction ........................................................................................................................................................................E.1.2
      • E.1.3
        The role of prefix length in our experiments ........................................................................................................................................................................E.1.3
    • E.2
      Setup ........................................................................................................................................................................E.2
    • E.3
      Results of varying prefix length on known training data ........................................................................................................................................................................E.3
    • E.4
      Results of negative control experiments ........................................................................................................................................................................E.4
      • E.4.1
        Results on Phi 4 ........................................................................................................................................................................E.4.1
      • E.4.2
        Results on books from after training cutoffs ........................................................................................................................................................................E.4.2
    • E.5
      Additional notes on “the monkey at the typewriter” metaphor ........................................................................................................................................................................E.5
    • E.6
      Additional notes on extraction and membership inference ........................................................................................................................................................................E.6
      • E.6.1
        The relationship between membership inference and extraction ........................................................................................................................................................................E.6.1
      • E.6.2
        We know ground-truth membership for most of our experiments ........................................................................................................................................................................E.6.2
      • E.6.3
        We are inferring membership for models with undisclosed training data ........................................................................................................................................................................E.6.3
      • E.6.4
        Our inferences about membership are about sequences, not whole books ........................................................................................................................................................................E.6.4
      • E.6.5
        Absence of extraction does not imply anything about membership ........................................................................................................................................................................E.6.5
      • E.6.6
        Extractability and copyright ........................................................................................................................................................................E.6.6
  • F
    Reconstructing books with one seed prompt ........................................................................................................................................................................F
    • F.1
      Reconstruction method ........................................................................................................................................................................F.1
    • F.2
      Running partial reconstruction for books with partial high memorization signal ........................................................................................................................................................................F.2
    • F.3
      Running reconstruction for books with low memorization signal ........................................................................................................................................................................F.3
  • G
    Extended discussion of results ........................................................................................................................................................................G
@@ A.1 Additional notes on the “careless people” example
The headline figure we include in Figure 1 (Section 2) is meant to be illustrative—an instance of, not a general statement about, memorization and extraction. This example also plays a role in some of our validity experiments (Section 5 and Appendix E). The text is drawn from The Great Gatsby [gatsby]; it is one of the most famous quotes from the book. It is also a short, $57$-token sequence: the prefix is $25$ tokens and the suffix is $32$ (with the Llama 1 tokenizer). We do not focus on such short suffixes in our analysis, for reasons that are discussed both in the main text and in our experimental details below. While extraction of verbatim $32$-token suffixes is often accepted as valid evidence for memorization [biderman2023pythia], it is now more common practice to use $50$-token suffixes. (Regarding validity, also refer to Section 5 and Appendix E.) We compute the extraction probability with respect to top-$k$ sampling ($T=1$, $k=40$). We do this by transforming the model logits according to the sampling parameters of this scheme. (See more about this below in Appendix A.3.)
It is also the case that very famous quotes like this are duplicated in many locations in training-data sources (e.g., blogs, essays, personal web pages, etc.) We are not claiming (here, or in general) that this sequence can be extracted because a given book (here, The Great Gatsby) was included in Books3; it could be extracted (or more extractable) because it is duplicated in many places in the training data. Of course, one of those copies is in Books3, if the entirety of Books3 is included in the model’s training data. We discuss this further in Appendix E.
@@ A.2 Metrics
The metric that we use for memorization comes from hayes2025measuringmemorizationlanguagemodels. The authors provide a different framing for the metric. Instead of directly comparing $p_{\bm{z}}$ values—for the same sequence across LLMs, across multiple sequences for the same LLM, etc.—hayes2025measuringmemorizationlanguagemodels specify a threat model with an adversary that can query the model an arbitrary number of times with the same prompt—an adversary that has $n$ independent attempts to extract a sequence $\bm{z}$} with at least probability $p$. This is where their metric derives its name, $\bm{(n,p)}$-discoverable extraction:
Definition 1 ($(n,p)$-discoverable extraction, from hayes2025measuringmemorizationlanguagemodels).
@@
On its own, this type of method is well-suited to completion models that continue the prompt text, rather than instruction-tuned chatbots that reply in a conversational style. We therefore only study completion models and, for brevity, refer to these models by name, version, and size. Also note that this definition can be (at least in principle) easily adapted to non-verbatim extraction by testing if a generation is within $\epsilon$ distance from the target suffix for a given distance metric.
Despite these differences in framing, in practice for verbatim extraction, computing this metric just involves calculating sequence probabilities $p_{{\bm{z}},\theta,\phi}$. For sequence $\bm{z}$}, this is the suffix probability with respect to an LLM’s base distribution $\theta$ and decoding scheme $\phi$ (e.g., greedy, top-$k$) that determines which tokens are considered at each step and how their logits are renormalized into conditional probabilities (Appendix A.3). That is,
@@
We will always note the particular LLM $\theta$ and decoding scheme $\phi$ being evaluated, and so, for brevity, will refer to this quantity as simply $p_{\bm{z}}$. In practice, we compute $p_{\bm{z}}$ as the $\exp$ of the sum of conditional $\log$ probabilities of the tokens in the sequence, as this is more numerically stable than multiplying together the conditional probabilities of the tokens in the sequence. (See Section 2.)
In Figure 10 (left), we show examples of $p_{\bm{z}}$ for the “careless people” quote (split into prefix and suffix as in Figure 1) for different models. (This is Figure 2 in Section 2, with Pythia 12B added.) Once one has computed $p_{\bm{z}}$, they can use $p_{\bm{z}}$ to compute $n$ for a given $p$ (for $p_{\bm{z}}\in(0,1)$) with
@@
Since $p_{\bm{z}}$ is the probability of generating the suffix with $1$ prompt to the model, the probability of not generating it in $1$ prompt is $(1-p_{\bm{z}})$. The probability of not generating it in $n$ independent prompts to the model (with the same prompt) is $(1-p_{\bm{z}})^{n}$, and so the probability of generating it in $n$ prompts is $1$ minus this probability, i.e., $1-(1-p_{\bm{z}})^{n}$.
As an intuition for this presentation of the metric, think about a fair coin flip coming up heads. Here, $p_{\bm{z}}=0.5$ (the probability of heads in one flip). With two actual flips ($n=2$), the probability of flipping heads at least once is $1-(1-0.5)^{2}=0.75=p$; when $n=10$, $p=0.9990$. We visualize this intuition in Figure 11 (left) with coin flips, and show how it translates to suffix generation probabilities for Llama 1 30B and the “careless people” quote, for which $p_{\bm{z}}=35.2\%$ (right). As is clear Equation (6), for a given $p_{\bm{z}}$, one can pick a probability threshold $p$ and get the corresponding number of prompts $n$ (or vice versa). Figure 10 (right) visualizes how $n$ changes for different settings of $p$ for the “careless people” quote.
In this paper, we re-frame $(n,p)$-discoverable extraction to emphasize the quantity $p_{\bm{z}}$ (A.2) that we actually compute: the probability of extracting a sequence $\bm{z}$} (for the given LLM, hyperparameter-configured decoding scheme, and suffix start index location within sequence $\bm{z}$}). This is also useful for plotting distributions over $p_{\bm{z}}$ for a given book—to see how probabilities vary for sequences across a book, or for a given sequence across different models. We provide such comparisons in Section 6.
In practice, for LLMs, we can compute $p_{\bm{z}}$ for verbatim extraction without prompting $n$ times. We compute sequence probabilities directly from the logits we obtain from running sequences through the LLM, rather than generating multiple suffixes for each prefix. (See Appendix A.3.) These two approaches are functionally equivalent for verbatim extraction: the computed sequence probabilities correspond to the statistically expected frequencies of verbatim extracted outputs we would observe if we were to generate a large number of sequences for the same prompt. hayes2025measuringmemorizationlanguagemodels, the authors of this extraction approach, confirm this with ample experimental evidence. We do, as well, though omit these results for brevity because they follow directly from what $p_{\bm{z}}$ is measuring.
We limit ourselves to verbatim extraction in this paper; it is significantly computationally cheaper to do so because we can operate directly on model logits with one forward pass. (See Section 2 and Appendix A.3 for more details.) As a result, our work here (in several respects) only scratches the surface: there are various different ways to instantiate this extraction methodology in practice, we only explore a limited set of open-weight models, and we only test $50$ of the nearly 200,000200{,}000 books in Books3.
We further discuss how we determine extraction success for probabilistic extraction in Appendix C, where we describe how extraction success for different sequences can be aggregated into an overall extraction rate. In the main text, we discuss determining extraction success in Section 2.1.